# Does DetectGPT Fully Utilize Perturbation? Selective Perturbation on Model-Based Contrastive Learning Detector would be Better

### 01 Feb 2024 v1

@@ 4.1 Experiment Settings
Few-shot Learning Settings. Experiments evaluating competitors and Pecola follow the setting of CoCo Liu et al. (2023). For the few-shot learning settings, we randomly sample 32, 64, 128 and 512 samples from the original training set, while keeping the balance of machine and human categories. More details are provided in Appendix A.
@@ 4.2 Comparison Models
GPT-2 Radford et al. (2019), RoBERTa Liu et al. (2019), supervised methods via standard fine-tuning pre-trained language models as classifiers. We use GPT-2 Small (124M) and RoBERTa-base (135M).
GLTR Gehrmann et al. (2019), an unsupervised MGT detection method based on next-token probability. We follow the setting of Guo et al. (2023), utilizing the Test-2 feature. For a fair comparison with fine-tuning methods, we first use the few-shot training samples to settle the threshold and adapt the fixed threshold in the test set.111The base model of GLTR is chosed based on the generator of the dataset: for GPT-2 and Grover datasets, we use GPT-2 Small (124M); and for GPT-3.5 and HC3 datasets, we use GPT-J (6B) Wang (2021), which is the best open-source model to simulate ChatGPT and GPT-3.5 empirically.
DetectGPT Mitchell et al. (2023), a curvature-based unsupervised zero-shot MGT detector, using T5-large Raffel et al. (2020) to perturb texts. Same as GLTR, we fix the threshold.222For all four datasets (including HC3 and GPT-3.5 datasets), we use GPT-2 Small (124M) as the base model to calculate the likelihood. The reason is Mireshghallah et al. (2023) find that small model is better black-box detector for DetectGPT.
@@
CE+SCL Gunel et al. (2021), a supervised contrastive learning method, used in conjunction with the Cross-Entropy (CE) loss, exhibiting impressive performance in few-shot learning settings.
CE+Margin Zhou et al. (2021), a contrastive learning approach focuses on separating Out-of-Distribution (OOD) instances from In-Distribution (ID) instances, aiming to minimize the L distance between instances of the same label. We train the detector by combining CE loss.
@@ 4.3 Performance Comparison
As shown in Table 1, Pecola surpasses the competitors on all datasets in the few-shot MGT detection task. Specifically, compared with the best competitor, Pecola achieves accuracy and F1-score improvement of 2.04% and 1.42%, 1.71% and 2.55% on Grover and GPT2 datasets. On GPT3.5 and HC3 datasets, Pecola still ensures 0.86% and 0.68%, 0.21% and 0.22% performance improvement with greater stability. The results prove the effectiveness of Pecola.
Moreover, the unsupervised learning methods tend to show better performance, and their performance improves with the increase in the number of training samples. Unsurprisingly, unsupervised methods do not see a notable performance improvement in this context, which causes them to outperform on the fewest shot settings initially but soon be surpassed.
Also, it should be noticed that the introduction of contrastive learning drastically improves the detector’s performance under a few-shot setting, demonstrating the necessity of utilizing both instance-level and class-level information from a few training samples. As for the deception of generators, Grover appears to be the hardest to detect while other models are relatively "honest" to detectors. It might have originated from the adversarial training strategy of Grover, while the bulit-in detector module adversarially shifts the LLM’s detectable features. More interestingly, advanced language models show a weaker ability to cheat detectors. Most detectors achieve around 98% in accuracy on the GPT-3.5 and HC3 datasets, which is consistent with the conclusion from Liu et al. (2023); Chen et al. (2023). We hypothesize that the easy-to-detect nature may originate from the lack of semantics diversity in GPT-3.5 and ChatGPT as they are using RLHF Kirk et al. (2023).
@@ 4.4 Ablation Study
To illustrate the effectiveness of the Pecola components, we study the ablation experiments on the Selective Strategy Perturbation stage and the Contrastive Learning stage on the 64-example GPT-2 dataset.
| GPT-2 | ||
@@
Ablation on Selective Strategy Perturbation. In the supervised MGT detection, the data used for training primarily includes original texts, selected mask texts, and mask-filled texts. We remove each part of the data in training, i.e., (i) w/o. mask, refers to not using selected mask texts for training; (ii) w/o. mask-fill, not using mask-filling texts for training.
Ablation on Contrastive Learning. It primarily investigates the impact of CE loss and contrastive loss. (i) w/o. CL refers to the model larking weighted contrastive learning; (ii) w/o. refers to the model including contrastive learning but larking weight.
As demonstrated in Table 2, in scenarios employing only the CE loss, the Selective Strategy Perturbation method contributes to significant performance improvement. Moreover, the introduction of weighting further enhances accuracy when compared to the direct use of margin loss. Furthermore, within the overall framework, the removal of the select mask text results in a more rapid decrease in accuracy compared to the removal of the mask-filling text. This finding substantiates that the Token-Level Weighted Multi-Pairwise Contrastive Learning method can better focus on the alterations in the rephrased space following the application of Selective Strategy Perturbation to the text.
4.5 Discussion
4.5.1 Model Qualities
We analyze the model qualities, including robustness, affinity, and diversity in this section. Here, we test on the 10,000-example GPT-2 test dataset, and the perturbation scale is set to 15%.
@@ 4.5.1 Model Qualities
Analysis on Affinity. Affinity pertains to alterations in data distribution resulting from perturbations, quantified by observing the fluctuations in accuracy. We demonstrate the superiority of the selective masking method over the random masking method using the Affinity metric, following the setting of DetectGPT Mitchell et al. (2023). We applied a 15% mask proportion with a span of 2 tokens on the test dataset and simultaneously employed T5-Large Raffel et al. (2020) as the mask-filling model. We trained RoBERTa-base and our Pecola on the 64-example GPT2 dataset. As shown in Table 4, in comparison to the random masking perturbation method utilized in DetectGPT, we observe a 1.92% and 0.49% increase in Affinity when employing the selective masking method. Additionally, the mask-filling method yields affinity improvements of 3.38% and 1.32% for RoBERTa and Pecola models, respectively. These results illustrate that the Selective Multi-Strategy Perturbation method introduced in this paper effectively preserves more distinguishable features between MGTs and HWTs.
@@
Analysis on Diversity. Conversely, diversity assesses the range and variability of perturbed data, utilizing metrics Dist-1 and Dist-2 Celikyilmaz et al. (2020). Here, we use three common perturbation methods to demonstrate the importance of not arbitrarily changing important tokens and the significance of select masks. (1) Token Substitution Zhang et al. (2015), replaces tokens with synonyms from WordNet Miller (1992); (2) SwitchOut Wang et al. (2018), uniformly samples and randomly substitutes from the vocabulary of test samples; and (3) Two-stage Wei et al. (2021) trains the mask-filling model on the original data.
| Method | Affinity | Diversity | |
| Dist-1 | Dist-2 | ||
| Original | - | 8.70 | 50.32 |
| Token Substitution | -20.00 | 3.38 | 43.43 |
| SwitchOut | -22.06 | 6.81 | 53.61 |
| Two-stage | -21.13 | 3.24 | 41.85 |
| Pecola Mask-Filling | -1.34 | 15.59 | 57.01 |
@@ 4.5.2 Generalization on Mask-Filling Models
We study the influence of various mask-filling models on the performance of Pecola, including Bert (110M; Devlin et al. 2019), Bart (139M; Mike et al. 2020), GPT-2 (380M; Radford et al. 2019), Twhin-bert (279M; Zhang et al. 2023), XLM (279M; Alexis et al. 2020), XLNet (110M; Yang et al. 2019), and RoBERTa(135M; Liu et al. 2019).
As depicted in Fig. 3, the result of every mask-filling model, in accuracy and F1-score, surpasses the baseline. Furthermore, the fluctuation of Pecola’s performance across different mask-filling models is relatively minor. It confirms that Pecola is not reliant on a specific filling model, showing great generalization capability. Detailed experiment settings and results are in Appendix C.
4.5.3 Effectiveness from domains and genres
We evaluate Pecola’s effectiveness from two aspects: domains and genres.
Cross-domain. We evaluate Pecola on the HC3 dataset, which primarily includes Medicine, Finance, and Computer Science domain QA. Each of the three domains is the training set respectively, while the remains from other domains are the test set. The results in Table 6 show that Pecola is more effectiveness than the baseline across all settings. For example, Pecola outperforms by 7.89% in Medicine, 3.26% in Finance, and 2.68% in Computer Science. It demonstrates the utility of selective perturbation and contrastive learning to maintain and even improve effectiveness.
| Domain | Medicine | Finance | Computer Science |
| RoBERTa | 62.97 | 86.08 | 90.64 |
| Pecola | 70.86 | 89.34 | 93.32 |
Cross-genre. We generalize Pecola between News articles (GPT3.5 dataset) and QA answers (HC3 dataset) on the 64-shot settings. In cross-genre, as shown in Table 7, where the GPT-3.5 dataset is the training set, Pecola outperforms by 14.19%; and where the HC3 dataset is the training set, Pecola outperforms by 10.2% to the baseline. It demonstrates our method’s advantage in the few-shot settings and also cross-genre.
| Dataset | GPT3.5 HC3 | HC3 GPT3.5 |
| RoBERTa | 64.60 | 62.67 |
| Pecola | 78.79 | 72.87 |
4.5.4 Detecting Shorter Texts
@@ 5 Conclusion
In this paper, we propose a novel MGT detection method, Pecola. To relieve the information loss caused by the random masking used in DetectGPT, we present a token-level selective strategy perturbation method. To more fully distinguish meaningful recombination spaces and reduce reliance on the mask-filling models, we present a token-level weighted multi-pairwise contrastive learning method. In few-shot learning settings, experimental results show that Pecola significantly enhances the performance of PLMs in MGT detection. Subsequent analytical experiments validate Pecola obtains good effectiveness, robustness, generalization, and capability in detecting short texts.
@@ Appendix A Implementation Details
This part mentions the implementation details and hyper-parameter settings of all the methods in the experiment as shown in Table 8. We randomly select 10 different seeds for experiments, and report average test accuracy and F1-score.
@@ B.2 Score Threshold
Appendix C Impact of the Chose mask-filling Models
This section shows the full experimental results of different mask-filling models, the experimental results confirm the same outcomes as in the few-shot learning settings, where the T5 filling model does not perform the best across all datasets. All the above models are obtained from huggingface transformers Wolf et al. (2020). And we do not intervene in the temperature sampling of the mask-filling model, setting it all to 1.

### 24 Feb 2024 v3

@@ 4.1 Experiment Settings
Few-shot Learning Settings. We randomly sample 32, 64, 128 and 512 samples from the original training set, while keeping the balance of machine and human categories. More details are provided in Appendix A.1.
@@ 4.2 Comparison Models
RoBERTa Liu et al. (2019), supervised methods via standard fine-tuning pre-trained language models as classifiers. We use RoBERTa-base (135M). GLTR Gehrmann et al. (2019), a metric-based detector and based on next-token probability. We follow the setting of Guo et al. (2023), utilizing the Test-2 feature. For a fair comparison with fine-tuning methods, we first use the few-shot training samples to settle the threshold and adapt the fixed threshold in the test set.111The base model of GLTR is chosen based on the generator of the dataset: for GPT-2 and Grover datasets, we use GPT-2 Small (124M); and for GPT-3.5 and HC3 datasets, we use GPT-J (6B) Wang (2021), which is the best open-source model to simulate ChatGPT and GPT-3.5 empirically.
CE+SCL Gunel et al. (2021), a fine-tuned detector, used in conjunction with the Cross-Entropy (CE) loss, exhibiting impressive performance in few-shot learning settings.
CE+Margin Zhou et al. (2021), a contrastive learning approach focuses on separating OOD instances from In-Distribution (ID) instances, aiming to minimize the L distance between instances of the same label. We train the detector by combining CE loss.
@@
DetectGPT Mitchell et al. (2023), a zero-shot metric-based MGT detector, using T5-large Raffel et al. (2020) to perturb texts. Same as GLTR, we fix the threshold.222For all four datasets (including HC3 and GPT-3.5 datasets), we use GPT-2 Small (124M) as the base model to calculate the likelihood. The reason is Mireshghallah et al. (2023) find that small model is better black-box detector for DetectGPT.
@@ 4.3 Performance Comparison
As shown in Table 1, Pecola surpasses the competitors on all datasets in the few-shot MGT detection task. Specifically, compared with the best competitor, Pecola achieves accuracy and F1-score improvement of 2.04% and 1.42%, 1.71% and 2.55% on Grover and GPT2 datasets. On GPT3.5 and HC3 datasets, Pecola still ensures 0.86% and 0.68%, 0.21% and 0.22% performance improvement with greater stability. The results prove the effectiveness of Pecola, which integrates the advantage of unsupervised (perturbation for metric-based) and supervised (contrastive learning for model-based) MGT detection methods.
Moreover, the unsupervised learning methods tend to show better performance in extremely few shot scenarios. Unsurprisingly, unsupervised methods do not see a notable performance improvement with the increase in the number of training samples, which causes them to outperform on the fewest shot settings initially but soon be surpassed. As for the deception of generators, Grover appears to be the hardest to detect while other models are relatively "honest" to detectors. It might have originated from the adversarial training strategy of Grover, while the bulit-in detector module adversarially shifts the LLM’s detectable features. More interestingly, advanced language models show a weaker ability to cheat detectors. Most detectors achieve around 98% in accuracy on the GPT-3.5 and HC3 datasets, which is consistent with the conclusion from Liu et al. (2023); Chen et al. (2023). We hypothesize that the easy-to-detect nature may originate from the lack of semantics diversity in GPT-3.5 and ChatGPT as they are using RLHF Kirk et al. (2023).
@@ 4.4 Ablation Study
To illustrate the effectiveness of the Pecola components, we study the ablation experiments on the Selective Strategy Perturbation stage and the Contrastive Learning stage on the 64-example GPT-2 dataset. We also demonstrate the Scalability of Pecola in Appendix C.
@@
Ablation on Selective Strategy Perturbation. In Pecola, the data used for training primarily includes original texts, selected mask texts, and mask-filled texts. We remove each part of the data in training, i.e., (i) w/o. mask, refers to not using selected mask texts for training; (ii) w/o. mask-fill, not using mask-filling texts for training.
Ablation on Contrastive Learning. It primarily investigates the impact of CE and contrastive loss. (i) w/o. CL refers to the model larking weighted contrastive learning; (ii) w/o. refers to the model including contrastive learning but larking weight.
As demonstrated in Table 2, in scenarios employing only the CE loss, the Selective Strategy Perturbation method contributes to significant performance improvement. Moreover, the introduction of weighting further enhances accuracy when compared to the direct use of margin loss. It reveals the validation of bridging the metric-based and model-based detectors, i.e., employing the Selective Strategy Perturbation method to evaluate the token importance for the multi-pairwise contrastive learning method. Furthermore, within the overall framework, the removal of the select mask text results in a more rapid decrease in accuracy compared to the removal of the mask-filling text. This finding substantiates that the Token-Level Weighted Multi-Pairwise Contrastive Learning method can better focus on the alterations in the rephrased space following the application of Selective Strategy Perturbation to the text.
@@
4.5 Discussion and Analysis
4.5.1 Model Qualities
We analyze the model qualities, including robustness and affinity in this section. Here, we test on the 10,000-example GPT-2 test dataset, and the perturbation scale is set to 15%.
Analysis on Affinity. Affinity pertains to alterations in data distribution resulting from perturbations, quantified by observing the fluctuations in accuracy. We demonstrate the superiority of the selective masking method over the random masking method using the Affinity metric, following the setting of DetectGPT. We applied a 15% mask proportion with a span of 2 tokens on the test dataset and simultaneously employed T5-Large Raffel et al. (2020) as the mask-filling model. We trained RoBERTa-base and Pecola on the 64-example GPT2 dataset. As shown in Table 4, in comparison to the random masking perturbation method utilized in DetectGPT, we observe a 1.92% and 0.49% increase in Affinity when employing the selective masking method. Additionally, the mask-filling method yields affinity improvements of 3.38% and 1.32% for RoBERTa and Pecola models, respectively. These results illustrate that the Selective Multi-Strategy Perturbation method introduced in this paper effectively preserves more distinguishable features between MGTs and HWTs.
@@ 4.5.1 Model Qualities
4.5.2 Analysis on Selective Strategies
Our method is a neat perturbation strategy that is mainly for the MGT detection task and empirically shows great performance. Beyond the Pecola and random perturbation method, we experiment with two other perturbation strategies: rank-based perturbation and keyword-based perturbation. In rank-based perturbation, we use the rescaled rank of next-token probability on GPT2-medium as the weight for perturbation position selection. In keyword-based perturbation, we prevent changes in the keywords extracted by the VLT-5 language model Pęzik et al. (2022) during perturbation. As shown in Table 5, the experimental results of selective perturbation outperform the random perturbation method by 1.20%, 2.04%, and 2.49% in average accuracy on the 64-example GPT2 dataset. The Importance-based strategy of Pecola is the highest.
| Method | Random | Prob. Rank | Keyword | Importance |
| Yake | 76.05 | 77.35 | 78.55 | 78.92 |
| Perplexity | 75.53 | 76.63 | 77.11 | 77.63 |
We further find that the random strategies of DetectGPT lead to more masking-filling failures than selective ones, which cause execution errors. Mask-filling failure ratio in Table 6 indicates that selective strategy based on token importance greatly decreases the failure ratio by 3.64%.
| Method | Random | Prob. Rank | Keyword | Importance |
| Ratio (%) | 9.20 | 7.83 | 7.80 | 5.56 |
@@ 4.5.3 Generalization on Mask-Filling Models
We study the influence of various mask-filling models on the performance of Pecola, including Bert (110M; Devlin et al. 2019), Bart (139M; Mike et al. 2020), GPT-2 (380M; Radford et al. 2019), Twhin-bert (279M; Zhang et al. 2023), XLM (279M; Alexis et al. 2020), XLNet (110M; Yang et al. 2019), RoBERTa(135M; Liu et al. 2019), and LLaMa-2(7B; Touvron et al. 2023). As depicted in Fig. 3, the results of all mask-filling models, in accuracy, surpass the baseline. Furthermore, the fluctuation of Pecola’s performance across different mask-filling models is relatively slight. It confirms that Pecola is not reliant on a specific filling model, showing great generalization capability. The experiment details and results are in Appendix D.1.
4.5.4 Generalization on Data
Cross-domain. We evaluate Pecola on the HC3 dataset, the meta-information details are in Appendix A.2. For the three domains of data, we use one of them as training data (64-shot), and the remaining domains of data as testing data. The results in Table 7 show that Pecola is more effective than the best baseline and SOTA method on average. For example, compared to Roberta, Pecola outperforms by 4.61% in three domains on average. And Pecola maintains a 1.63% higher accuracy on average than SOTA DetectGPT.
| Domain | Medicine | Finance | Computer Sci. | Average |
| RoBERTa | 62.97 | 86.08 | 90.64 | 79.90 |
| DetectGPT | 80.48 | 85.17 | 82.98 | 82.88 |
| Pecola | 70.86 | 89.34 | 93.32 | 84.51 |
Cross-genre. We generalize Pecola between News articles (GPT3.5 dataset) and QA answers (HC3 dataset) on the 64-shot settings. As shown in Table 8, when the GPT-3.5 dataset is the training set, Pecola outperforms by 10.21%; and when the HC3 dataset is the training set, Pecola outperforms by 6.98% to the best competitor.
| Dataset | GPT3.5HC3 | HC3GPT3.5 | Average |
| RoBERTa | 64.60 | 62.67 | 63.64 |
| DetectGPT | 77.11 | 72.66 | 74.89 |
| Pecola | 78.79 | 72.87 | 75.83 |
@@ 5 Conclusion
In this paper, we introduce Pecola, a novel machine-generated text detection method that effectively bridges and integrates metric-based and fine-tuned detectors for MGT detection. To relieve the information loss caused by the random masking used in DetectGPT, we present a token-level selective strategy perturbation method. To more fully distinguish meaningful recombination spaces and reduce reliance on the mask-filling models, we present a token-level weighted multi-pairwise contrastive learning method. In few-shot settings, experimental results show that Pecola significantly enhances the performance of PLMs in MGT detection. Subsequent analytical experiments validate Pecola’s effectiveness, robustness, generalization, and capability in detecting short texts.
@@ Appendix A Implementation Details
This part mentions the hyperparameter settings and meta-information of the HC3 dataset.
A.1 Hyperparameter Details
Experiments evaluating competitors and Pecola follow the setting of CoCo Liu et al. (2023). The hyperparameter settings of all the methods in the experiment as shown in Table 9. We randomly select 10 different seeds for experiments, and report average test accuracy and F1-score.
@@ A.1 Hyperparameter Details
A.2 Meta-information
We evaluate Pecola effectiveness from domains and genres on the HC3 dataset, which primarily includes Medicine, Finance, and Computer Science domain QA, as shown in Table 10.
| Domain | Medicine | Finance | Computer Science |
| Size | 2585 | 8436 | 1684 |
@@ B.2 Score Threshold
Appendix C Efficiency and Scalability of Pecola
We adopt Pythia Biderman et al. (2023) as the base model of Pecola with different scales, i.e., 70M, 160M, 410M, 1B, and 1.4B. We train and do experiments on one NVIDIA A100 GPU, and the performance and time consumption are in Table 11. With the increase in model size, both accuracy and F1-score show upward trends, while the time increase is linear, which is reasonable.
| Model | 70M | 160M | 410M | 1B | 1.4B |
| Acc | 58.42 | 63.66 | 71.07 | 72.13 | 74.05 |
| F1 | 58.03 | 63.54 | 70.87 | 71.75 | 73.85 |
| Per epoch | 16s | 34s | 85s | 97s | 113s |
| Single data | 2.2ms | 7.0ms | 13.8ms | 14.1ms | 16.6ms |
Appendix D Further Clarification on Perturbation
Conversely, diversity assesses the range and variability of perturbed data, utilizing metrics Dist-1 and Dist-2 Celikyilmaz et al. (2020). Here, we use three common perturbation methods to demonstrate the importance of not arbitrarily changing important tokens and the significance of select masks. (1) Token Substitution Zhang et al. (2015), replaces tokens with synonyms from WordNet Miller (1992); (2) SwitchOut Wang et al. (2018), uniformly samples and randomly substitutes from the vocabulary of test samples; and (3) Two-stage Wei et al. (2021) trains the mask-filling model on the original data.
| Method | Affinity | Diversity | |
| Dist-1 | Dist-2 | ||
| Original | - | 8.70 | 50.32 |
| Token Substitution | -20.00 | 3.38 | 43.43 |
| SwitchOut | -22.06 | 6.81 | 53.61 |
| Two-stage | -21.13 | 3.24 | 41.85 |
| Pecola Mask-Filling | -1.34 | 15.59 | 57.01 |
D.1 Impact of the Chosen Mask-filling Models
This section shows the full experimental results of different mask-filling models, as shown in Table 13, the experimental results confirm the same outcomes as in the few-shot learning settings, where the T5 filling model does not perform the best across all datasets. All the above models are obtained from huggingface transformers Wolf et al. (2020). And we do not intervene in the temperature sampling of the mask-filling model, setting it all to 1.